Maria and Juan have some m&m's. Since Maria has three times as juan, he started crying. So Maria decided to give him 3 m&m's. She now has twice as much as Juan. How many m&m's does each person have now?
Answer by Theo(13342)

x is the amount the juan has.
3x is the amount that maria has.

maria gives juan 3 m&m.

she now has 3x - 3

he now has x + 3

shw now has 2 times as many m^m as he has.

3x-3 = 2*(x+3)

simplify this equation to get:

3x-3 = 2x+6

subtract 2x from both sides of this equation and add 3 to both sides of this equation to get:

x = 9

x is equal to the amount that juan has.

juan has 9
maria has 27.

maria has 3 times as many as juan.

maria gives juan 3.

maria has 24
juan has 12

maria now has 2 times as many as juan has.

formula is good.

solution is:

maria now has 24
juan now has 12